# Quantum Algorithms for the Pathwise Lasso

### v2

@@ Abstract
We present a novel quantum high-dimensional linear regression algorithm with an $\ell_{1}$-penalty based on the classical LARS (Least Angle Regression) pathwise algorithm. Similarly to available classical numerical algorithms for Lasso, our quantum algorithm provides the full regularisation path as the penalty term varies, but quadratically faster per iteration under specific conditions. A quadratic speedup on the number of features/predictors $d$ is possible by using the simple quantum minimum-finding subroutine from Dürr and Høyer (arXiv’96) in order to obtain the joining time at each iteration. We then improve upon this simple quantum algorithm and obtain a quadratic speedup both in the number of features $d$ and the number of observations $n$ by using the approximate quantum minimum-finding subroutine from Chen and de Wolf (ICALP’23). In order to do so, we construct, as one of our main contributions, a quantum unitary based on quantum amplitude estimation to approximately compute the joining times to be searched over by the approximate quantum minimum-finding subroutine. Since the joining times are no longer exactly computed, it is no longer clear that the resulting approximate quantum algorithm obtains a good solution. As our second main contribution, we prove, via an approximate version of the KKT conditions and a duality gap, that the LARS algorithm (and therefore our quantum algorithm) is robust to errors. This means that it still outputs a path that minimises the Lasso cost function up to a small error if the joining times are only approximately computed. Furthermore, we show that, when the observations are sampled from a Gaussian distribution, our quantum algorithm’s complexity only depends polylogarithmically on $n$, exponentially better than the classical LARS algorithm, while keeping the quadratic improvement on $d$. Finally, we propose a dequantised version of our quantum algorithm that also retains the polylogarithmic dependence on $n$, albeit presenting the linear scaling on $d$ from the standard LARS algorithm.
@@ 1 Introduction
    As $\ell_{1}$-penalties are convex, sparsity leads to computational advantages. For example, take two million features and only 200 observations, then estimating two million parameters becomes extremely difficult.
@@ 1.1 Least Angle Regression algorithm
$\hat{\bm{\beta}}(\lambda)\in\underset{\bm{\beta}\in\mathbb{R}^{d}}{\mathrm{argmin}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{for all}\leavevmode\nobreak\ \lambda>0.$ (1)
@@
There is a maximal value $\lambda_{\max}=\lambda_{0}$ where $\hat{\bm{\beta}}(\lambda)=0$ for all $\lambda\geq\lambda_{0}$ and $\hat{\beta}_{j}(\lambda)\neq 0$ for some $j\in[d]$ and $\lambda<\lambda_{0}$. Such value is $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$. The $m+1$ points $\lambda_{0},\dots,\lambda_{m}$ where $\partial\hat{\bm{\beta}}(\lambda)/\partial\lambda$ changes are called kinks, and the path $\{\hat{\bm{\beta}}(\lambda):\lambda_{t+1}<\lambda\leq\lambda_{t}\}$ between two consecutive kinks $\lambda_{t+1}$ and $\lambda_{t}$ defines a linear segment. The fact that the regularisation path is piecewise linear and the locations of the kinks can be derived from the Karush-Kuhn-Tucker (KKT) optimality conditions (see Section 3 for more details).
@@ 1.2 Our work
Given $\epsilon:\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$, we say that a vector $\widetilde{\bm{\beta}}\in\mathbb{R}^{d}$ is an approximate Lasso solution with error $\epsilon(\lambda)$ if
$\frac{1}{2}\|\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}\|_{2}^{2}+\lambda\|\widetilde{\bm{\beta}}\|_{1}-\left(\min_{\bm{\beta}\in\mathbb{R}^{d}}\frac{1}{2}\|\mathbf{y}-\mathbf{X}\bm{\beta}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\right)\leq\epsilon(\lambda)\quad\text{where}\leavevmode\nobreak\ \lambda>0.$
@@
Result 2 (Informal version of Theorem 28).
@@
Result 2 guarantees the correctness of all our algorithms. Their time complexities, on the other hand, are given by the theorem below and summarised in Table 1.
Result 3 (Informal version of Theorem 25, Theorem 29, Theorem 30).
@@
Classical LARS 1 None $O(n|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2})$
Simple quantum LARS 2 None $\widetilde{O}(n\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2})$
Approximate classical LARS 4 $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}\left(\frac{\gamma_{\mathcal{A}}^{-2}+n\|\mathbf{X}\|_{\max}^{2}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}^{2}}{(1-\alpha_{\mathcal{A}})^{2}\epsilon^{2}}|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
Approximate quantum LARS 3 $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}\left(\frac{\gamma_{\mathcal{A}}^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}}{(1-\alpha_{\mathcal{A}})\epsilon}\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2}\right)$
Table 1: Summary of results. Throughout this work, $n$ is the number of observations, $d$ is the number of features, $\mathcal{A}$ is the active set, and $\mathcal{I}$ is the inactive set. In addition, $\epsilon>0$ is the approximation error, $\delta\in(0,1)$ is the failure probability, $T$ is the number of kinks in the regularisation path $\mathcal{P}$, $\mathbf{X}_{\mathcal{A}}=[\mathbf{X}_{i}]_{i\in\mathcal{A}}\in\mathbb{R}^{n\times|\mathcal{A}|}$ is the matrix formed by the columns of $\mathbf{X}$ in $\mathcal{A}$, $\mathbf{X}^{+}_{\mathcal{A}}$ is the Moore–Penrose inverse of $\mathbf{X}_{\mathcal{A}}$, and the parameters $\alpha_{\mathcal{A}},\gamma_{\mathcal{A}}\in(0,1)$ are defined in Result 3. The notation $\widetilde{O}(\cdot)$ omits $\operatorname{poly}\log$ terms in $n$, $d$, $T$, and $\delta$.
@@
Result 4 (Informal version of Lemma 31, Lemma 32, Lemma 33).
@@
Result 4 implies that the complexity of our approximate LARS algorithms, both quantum and classical, can be substantially better than the standard LARS algorithm depending on the input data. In the case when $\mathbf{X}$ is a Gaussian matrix and $\mathbf{y}$ is sparse, so that $\|\mathbf{y}\|_{1}/\|\mathbf{y}\|_{2}=\operatorname{poly}\log{n}$, the approximate quantum and classical complexities per iteration are $\widetilde{O}(\sqrt{d}/\epsilon)$ and $\widetilde{O}(d/\epsilon^{2})$, respectively, which only depend polylogarithmically on the number of samples $n$ (assuming that $n|\mathcal{A}|=O(\sqrt{d})$). The quantum runtime particularly is more than quadratically better than the usual $O(nd)$ classical runtime (at the expense of approximating the solution). The condition that $\mathbf{y}$ is sparse can be artificially generated through setting to zero some of the regression coefficients of the underlying true solution $\bm{\beta}^{\ast}$ connecting $\mathbf{X}$ and $\mathbf{y}$ (see Section 5). We note that sparse inputs have also provided speedups in other settings [6, 46, 7, 82, 64, 44, 62]. Independently, we believe that this result would be of interest to the compressed sensing community. The complexities for the case of random Gaussian design matrix $\mathbf{X}$ are summarised in Table 2.
@@
Classical LARS 1 None $O(nd)$
Simple quantum LARS 2 None $\widetilde{O}(n\sqrt{d}+n|\mathcal{A}|)$
@@
Our approximate LARS algorithms output a regularisation path with point-wise error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ (see Definition 1). Even though this is not a purely additive error $\epsilon$, it is still better than a purely multiplicative error $\epsilon\big{(}\frac{1}{2}\|\mathbf{y}-\mathbf{X}\widetilde{\bm{\beta}}\|_{2}^{2}+\lambda\|\widetilde{\bm{\beta}}\|_{1}\big{)}$. Moreover, this kind of error is standard in the literature, see Section 5 and also [9, 78, 114].
Compared to the previous work of Chen and de Wolf [23] on quantum algorithms for Lasso, our work is different in that we study the pathwise Lasso regression problem as the tuning parameter $\lambda$ varies, and not focus on a single point $\lambda$. It looks infeasible to retrieve even a small interval $[\hat{\bm{\beta}}(\lambda),\hat{\bm{\beta}}(\lambda^{\prime})]$ of Lasso solutions via their algorithms. Both works are thus incomparable in terms of results. The output solution from the quantum algorithm of Chen and de Wolf can actually be used to initialise our algorithms instead of the point $(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0})$. Regarding quantum techniques, on the other hand, both works employ similar quantum subroutines like quantum amplitude estimation and quantum minimum finding. The fact that no advanced subroutines are needed to improve the dependence on both $n$ and $d$ should be seen as a strength of our (and [23]) work. Finally, we briefly mention that the lower bound $\Omega(\sqrt{d}/\epsilon^{1.5})$ from [23] does not contradict our quantum upper bound $\widetilde{O}(\sqrt{d}/\epsilon)$ since Chen and de Wolf assume an additive error $\epsilon$.
@@ 1.3 Related work
While quantum linear regression has been relatively well studied in the unregularised setting, research on algorithms for regression using regularised least squares is not yet well established. In the sparse access model, Yu et al. [123] presented a quantum algorithm for Ridge regression using a technique called parallel Hamiltonian simulation to develop the quantum analogue of $K$-fold cross-validation [47], which serves as an efficient estimator of Ridge regression. Other works on quantum Ridge regression algorithms in the sparse access model include [102, 96, 24]. Chen and de Wolf [23] designed quantum algorithms for Lasso and Ridge regression from the perspective of empirical loss minimisation. Both of their quantum algorithms output a classical vector whose loss is $\epsilon$-close to the minimum achievable loss. Around the same time, Bellante and Zanero [14] gave a polynomial speedup for the classical matching-pursuit algorithm, which is a heuristic algorithm for the best subset selection model. More recently, Chakraborty et al. [17] gave the first quantum algorithms for least squares with general $\ell_{2}$-norm regularisation, which includes regularised versions of quantum ordinary least squares, quantum weighted least squares, and quantum generalised least squares. Their algorithms use block-encoding techniques and the framework of quantum singular value transformation [44], and demonstrate substantial improvement compared to previous results on quantum Ridge regression [102, 24, 123].

### v3

@@ Abstract
We present a novel quantum high-dimensional linear regression algorithm with an $\ell_{1}$-penalty based on the classical LARS (Least Angle Regression) pathwise algorithm. Similarly to available classical numerical algorithms for Lasso, our quantum algorithm provides the full regularisation path as the penalty term varies, but quadratically faster per iteration under specific conditions. A quadratic speedup on the number of features/predictors $d$ is possible by using the simple quantum minimum-finding subroutine from Dürr and Høyer (arXiv’96) in order to obtain the joining time at each iteration. We then improve upon this simple quantum algorithm and obtain a quadratic speedup both in the number of features $d$ and the number of observations $n$ by using the approximate quantum minimum-finding subroutine from Chen and de Wolf (ICALP’23). In order to do so, we construct a quantum unitary based on quantum amplitude estimation to approximately compute the joining times to be searched over by the approximate quantum minimum-finding subroutine. Since the joining times are no longer exactly computed, it is no longer clear that the resulting approximate quantum algorithm obtains a good solution. As another main contribution, we prove, via an approximate version of the KKT conditions and a duality gap, that the LARS algorithm (and therefore our quantum algorithm) is robust to errors. This means that it still outputs a path that minimises the Lasso cost function up to a small error if the joining times are only approximately computed. Furthermore, we show that, when the observations are sampled from a Gaussian distribution, our quantum algorithm’s complexity only depends polylogarithmically on $n$, exponentially better than the classical LARS algorithm, while keeping the quadratic improvement on $d$. Moreover, we propose a dequantised version of our quantum algorithm that also retains the polylogarithmic dependence on $n$, albeit presenting the linear scaling on $d$ from the standard LARS algorithm. Finally, we prove query lower bounds for classical and quantum Lasso algorithms.
@@ 1 Introduction
    As $\ell_{1}$-penalties are convex, sparsity leads to computational advantages. For example, if we have two million features but only 200 observations, estimating two million parameters becomes extremely difficult.
@@ 1.1 Least Angle Regression algorithm
$\hat{\bm{\beta}}\in\operatorname*{arg\,min}_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta})\quad\text{where}\quad\mathcal{L}^{(\lambda)}(\bm{\beta})\triangleq\frac{1}{2}\|\mathbf{y}-\mathbf{X}{\bm{\beta}}\|_{2}^{2}+\lambda\|\bm{\beta}\|_{1}\quad\text{for all}\leavevmode\nobreak\ \lambda>0.$ (1)
@@
There is a maximal value $\lambda_{\max}=\lambda_{0}$ where $\hat{\bm{\beta}}(\lambda)=\mathbf{0}$ for all $\lambda\geq\lambda_{0}$ and $\hat{\beta}_{j}(\lambda)\neq 0$ for some $j\in[d]$ and $\lambda<\lambda_{0}$. Such value is $\lambda_{0}=\|\mathbf{X}^{\top}\mathbf{y}\|_{\infty}$. The $m+1$ points $\lambda_{0},\dots,\lambda_{m}$ where $\partial\hat{\bm{\beta}}(\lambda)/\partial\lambda$ changes are called kinks, and the path $\{\hat{\bm{\beta}}(\lambda):\lambda_{t+1}<\lambda\leq\lambda_{t}\}$ between two consecutive kinks $\lambda_{t+1}$ and $\lambda_{t}$ defines a linear segment. The fact that the regularisation path is piecewise linear and the locations of the kinks can be derived from the Karush-Kuhn-Tucker (KKT) optimality conditions (see Section 3 for more details).
@@ 1.2 Our work
Given $\epsilon:\mathbb{R}_{>0}\to\mathbb{R}_{\geq 0}$, we say that a vector $\widetilde{\bm{\beta}}\in\mathbb{R}^{d}$ is an approximate Lasso solution with error $\epsilon(\lambda)$, or an $\epsilon(\lambda)$-minimiser, if
$\mathcal{L}^{(\lambda)}(\widetilde{\bm{\beta}})-\min_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta})\leq\epsilon(\lambda),\quad\text{where}\quad\lambda>0.$
@@
Result 1 (Informal version of Theorem 26).
@@
Result 1 guarantees the correctness of all our algorithms. Their time complexities, on the other hand, are given by the theorem below and summarised in Table 1.
Result 2 (Informal version of Theorems 22, 27 and 28).
@@
Classical LARS 1 00 $O(n|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2})$
Simple quantum LARS 2 00 $\widetilde{O}(n\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2})$
Approximate classical LARS 4 $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}\bigg{(}\frac{\gamma_{\mathcal{A}}^{-2}+n\|\mathbf{X}\|_{\max}^{2}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}^{2}}{(1-\alpha_{\mathcal{A}})^{2}\epsilon^{2}}|\mathcal{I}|+n|\mathcal{A}|+|\mathcal{A}|^{2}\bigg{)}$
Approximate quantum LARS 3 $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ $\widetilde{O}\bigg{(}\frac{\gamma_{\mathcal{A}}^{-1}+\sqrt{n}\|\mathbf{X}\|_{\max}\|\mathbf{X}^{+}_{\mathcal{A}}\|_{2}}{(1-\alpha_{\mathcal{A}})\epsilon}\sqrt{|\mathcal{I}|}+n|\mathcal{A}|+|\mathcal{A}|^{2}\bigg{)}$
Table 1: Summary of results. Throughout this work, $n$ is the number of observations, $d$ is the number of features, $\mathcal{A}$ is the active set, and $\mathcal{I}$ is the inactive set. In addition, $\epsilon>0$ is the approximation error, $\delta\in(0,1)$ is the failure probability, $T$ is the number of kinks in the regularisation path $\mathcal{P}$, $\mathbf{X}_{\mathcal{A}}=[\mathbf{X}_{i}]_{i\in\mathcal{A}}\in\mathbb{R}^{n\times|\mathcal{A}|}$ is the matrix formed by the columns of $\mathbf{X}$ in $\mathcal{A}$, $\mathbf{X}^{+}_{\mathcal{A}}$ is the Moore–Penrose inverse of $\mathbf{X}_{\mathcal{A}}$, and the parameters $\alpha_{\mathcal{A}},\gamma_{\mathcal{A}}\in(0,1)$ are defined in Result 2. The notation $\widetilde{O}(\cdot)$ omits $\operatorname{poly}\log$ terms in $n$, $d$, $T$, and $\delta$.
Our approximate LARS algorithms output a regularisation path with point-wise error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$. Even though this is not a purely additive error $\epsilon$, it is still better than a purely multiplicative error $\epsilon\mathcal{L}^{(\lambda)}(\widetilde{\bm{\beta}})$. It is possible to prove that $\mathcal{L}^{(\lambda)}(\widetilde{\bm{\beta}})\leq\frac{1}{1-\epsilon}\min_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta})$ (see Remark 25), therefore our approximate LARS algorithms output a solution with error at most $\frac{\epsilon}{1-\epsilon}\min_{\bm{\beta}\in\mathbb{R}^{d}}\mathcal{L}^{(\lambda)}(\bm{\beta})$. Moreover, this type of error $\lambda\epsilon\|\widetilde{\bm{\beta}}\|_{1}$ is standard in the literature, see Section 6 and also [10, 80, 116].
@@
Result 3 (Informal version of Lemma 29, 30 and 31).
@@
Result 3 implies that the complexity of our approximate LARS algorithms, both quantum and classical, can be substantially better than the standard LARS algorithm depending on the input data. In the case when $\mathbf{X}$ is a Gaussian matrix and $\mathbf{y}$ is sparse, so that $\|\mathbf{y}\|_{1}/\|\mathbf{y}\|_{2}=\operatorname{poly}\log{n}$, the approximate quantum and classical complexities per iteration are $\widetilde{O}(\sqrt{d}/\epsilon)$ and $\widetilde{O}(d/\epsilon^{2})$, respectively, which only depend polylogarithmically on the number of samples $n$ (assuming that $n|\mathcal{A}|=O(\sqrt{d})$). The quantum runtime, in particular, is more than quadratically better than the usual $O(nd)$ classical runtime (at the expense of approximating the solution). The condition that $\mathbf{y}$ is sparse can be artificially generated through setting to zero some of the regression coefficients of the underlying true solution $\bm{\beta}^{\ast}$ connecting $\mathbf{X}$ and $\mathbf{y}$ (see Section 6). We note that sparse inputs have also provided speedups in other settings [6, 47, 7, 84, 65, 45, 63]. Independently, we believe that this result would be of interest to the compressed sensing community. The complexities for the case of random Gaussian design matrix $\mathbf{X}$ are summarised in Table 2.
Finally, we prove query lower bounds for any classical and quantum algorithms for Lasso. The proof techniques come mostly from the work of Chen and de Wolf [24] but adapted to the Lasso problem from Equation 1 instead of $\min_{\bm{\beta}\in\mathbb{R}^{n}}\frac{1}{n}\|\mathbf{y}-\mathbf{X}\bm{\beta}\|_{2}^{2}$ such that $\|\bm{\beta}\|_{1}\leq 1$.
Result 4 (Theorems 43 and 44).
Let $n<d$ such that $n=\Omega(\log{d})$, $\lambda=\Omega(n)$, and $\epsilon\in(0,1/5)$ such that $\epsilon=\Omega(\sqrt{n^{3}\log{d}}/\lambda^{2})$. There is $\mathbf{X}\in\{-1,1\}^{n\times d}$ such that every bounded-error classical or quantum algorithm that computes an $\epsilon\lambda\|\widetilde{\bm{\beta}}\|_{1}$-minimiser for the Lasso $\mathcal{L}^{(\lambda)}(\bm{\beta})$ uses
  • classical: $\Omega\Big{(}\frac{n^{4}d}{\lambda^{4}\epsilon^{2}\log{d}}\Big{)}$ queries to $\mathbf{X}$. For $\epsilon=\Theta(\sqrt{n^{3}\log{d}}/\lambda^{2})$, this is $\Omega\Big{(}\frac{nd}{\log^{2}{d}}\Big{)}$;
  • quantum: $\Omega\Big{(}\frac{n^{3}\sqrt{d}}{\lambda^{3}\epsilon^{3/2}}\Big{)}$ queries to $\mathbf{X}$. For $\epsilon=\Theta(\sqrt{n^{3}\log{d}}/\lambda^{2})$, this is $\Omega\Big{(}\frac{n^{3/4}\sqrt{d}}{\log^{3/4}{d}}\Big{)}$.
We stress that our lower bound does not contradict the results from Table 2 since it is a worst-case bound: there is a design matrix $\mathbf{X}$ such that any classical or quantum algorithm requires $\Omega\Big{(}\frac{n^{4}d}{\lambda^{4}\epsilon^{2}\log{d}}\Big{)}$ or $\Omega\Big{(}\frac{n^{3}\sqrt{d}}{\lambda^{3}\epsilon^{3/2}}\Big{)}$ queries to $\mathbf{X}$, respectively, which does not exclude the existence of easy instances, a trivial case being the all-00 matrix. Still, for the range of parameters $\lambda=\Theta(n)$ and $\epsilon=\Theta(\sqrt{n^{3}\log{d}}/\lambda^{2})=\Theta(1)$, the time complexity per iteration of our approximate quantum LARS algorithm with a random Gaussian matrix is $\widetilde{O}(\sqrt{d})$, while for the approximate classical LARS it is $\widetilde{O}(d)$. Together with $O(n)$ iterations in order to reach the desired $\lambda$, their final complexities would be $\widetilde{O}(nd)$ (classical) and $\widetilde{O}(n\sqrt{d})$ (quantum), larger than their respective lower bounds $\widetilde{\Omega}(nd)$ and $\widetilde{\Omega}(n^{3/4}\sqrt{d})$. We note that, in this case, the time complexity of the usual classical LARS algorithm is $O(n^{2}d)$ (although the solution has no error).
@@
Classical LARS 1 00 $O(nd)$
Simple quantum LARS 2 00 $\widetilde{O}(n\sqrt{d}+n|\mathcal{A}|)$
@@
Compared to the previous work of Chen and de Wolf [24] on quantum algorithms for Lasso, our work is different in that we study the pathwise Lasso regression problem as the tuning parameter $\lambda$ varies, and not focus on a single point $\lambda$. It looks infeasible to retrieve even a small interval $[\hat{\bm{\beta}}(\lambda),\hat{\bm{\beta}}(\lambda^{\prime})]$ of Lasso solutions via their algorithms. On the other hand, if one wishes to compute $\hat{\bm{\beta}}(\lambda)$ for a specific $\lambda$, then their algorithms would be preferable. Both works are thus incomparable in terms of results. The output solution from the quantum algorithm of Chen and de Wolf can actually be used to initialise our algorithms instead of the point $(\lambda_{0},\hat{\bm{\beta}}(\lambda_{0})=\mathbf{0})$. Regarding quantum techniques, both works employ similar quantum subroutines like quantum amplitude estimation and quantum minimum finding. The fact that no advanced subroutines are needed to improve the dependence on both $n$ and $d$ should be seen as a strength of our (and [24]) work.
@@ 1.3 Related work
While quantum linear regression has been relatively well studied in the unregularised setting, research on algorithms for regression using regularised least squares is not yet well established. In the sparse access model, Yu et al. [125] presented a quantum algorithm for Ridge regression using a technique called parallel Hamiltonian simulation to develop the quantum analogue of $K$-fold cross-validation [48], which serves as an efficient estimator of Ridge regression. Other works on quantum Ridge regression algorithms in the sparse access model include [104, 98, 25]. Chen and de Wolf [24] designed quantum algorithms for Lasso and Ridge regression from the perspective of empirical loss minimisation. Both of their quantum algorithms output a classical vector whose loss is $\epsilon$-close to the minimum achievable loss. Around the same time, Bellante and Zanero [15, 8] gave a polynomial speedup for the classical matching-pursuit algorithm, which is a heuristic algorithm for the best subset selection model. More recently, Chakraborty et al. [18] gave the first quantum algorithms for least squares with general $\ell_{2}$-norm regularisation, which includes regularised versions of quantum ordinary least squares, quantum weighted least squares, and quantum generalised least squares. Their algorithms use block-encoding techniques and the framework of quantum singular value transformation [45], and demonstrate substantial improvement compared to previous results on quantum Ridge regression [104, 25, 125].